How many mg of metal containing 45% nickel must be combined with 6 mg of pure nickel to form an alloy containing78% nickel?
Answer by stanbon(75887)

How many mg of metal containing 45% nickel must be combined with 6 mg of pure nickel to form an alloy containing78% nickel?
-------
Equation:
nick + nick = nick
0.45x + 1*6 = 0.78(x+6)
------
45x + 100*6 = 78x + 78*6
------
-33x = -22*6
----
x = (2/3)6
x = 4 mg (amt. of pure nickel needed)
=====
Cheers,
Stan H.